OpenStudy (anonymous):

Tell if a triangle can have side lengths of 7, 9, and 11. If so, classify the triangle as acute, obtuse, or right. I know that it can make a triangle, however I am not sure on what type of triangle. I cannot remember how to figure that out. Please help=)

ganeshie8 (ganeshie8):

based on side lengths we can figure out if its scalene/isosceles/equilateral. but how to figure out if its acute/obtuse or right hmm ? looks tricky..

ganeshie8 (ganeshie8):

OK, here it is :- pick the longest side

ganeshie8 (ganeshie8):

square it and see if it equals sum of squares of other two sides

ganeshie8 (ganeshie8):

7, 9, 11 longest side = 11^2 = 121 sum of squares of other two sides = 7^2 + 9^2 = 130

ganeshie8 (ganeshie8):

since 121 < 130 that is, 11^2 < 7^2 + 9^2, the triangle is an acute triangle

OpenStudy (anonymous):

Thank you=)

ganeshie8 (ganeshie8):

if you get it >, then it wud be obtuse, if you get it =, then it wud be right. since we got <, in our present problem, it is acute.

ganeshie8 (ganeshie8):

np, yw :)